This quarter, there is a 50% chance that Jason will pass accounting, a 60% chance the he will pass english, and 80% chance that he will pass at least one of these two courses. What is the probability that he will pass both accounting and english?
Found 3 solutions by stanbon, ikleyn, greenestamps:
Answer by stanbon(75887)

This quarter, there is a 50% chance that Jason will pass accounting, a 60% chance the he will pass english, and 80% chance that he will pass at least one of these two courses. What is the probability that he will pass both accounting and english?
------
P(pass account and eng) = 0.5*0.6 = 0.3
------------
Cheers,
Stan H.
---------

Answer by ikleyn(52784)

.
Use the general formula of the Probability theory 


    P(Accounting OR English)) = P(Acc) + P(Eng) - P(Acc AND English).  


With the given data, this equation / (equality) takes the form


      80%                     = 50%    +  60%   - P(Acc AND Eng)   ====>  (it implies that)

  
    P(Acc AND Eng) = 50% + 60% - 80% = 30%.


Answer.  The probability that he will pass both Accounting and English is 30%, under the given conditions.


Answer by greenestamps(13200)



The problem is over-constrained; however, the constraints are consistent with one another.

To calculate the probability that he passes both classes, you only need to multiply the probabilities of his passing each class: 0.5*0.6 = 0.3.

With those numbers, the given information that there is an 80% chance of passing at least one of the courses is extraneous.

P(pass both) = 0.3
P(pass accounting only) = 0.5-0.3 = 0.2
P(pass english only) = 0.6-0.3 = 0.3

It follows (without having to be specified in the problem) that
P(pass at least one) = 0.3+0.2+0.3 = 0.8.

So the more complicated calculation shown by the other tutor, while valid, is more work than is required.